OpenStudy (anonymous):

Multiply. x^2 + 8x +15/x - 4 times x^2 -16/2x +6 A.x^4+8x-1/3x + 2 B.(x + 5)(x +4)/2 C.(x - 5) (x+4)/ 2 D.x^4+8x^3-x/ 2x^3 - 24

OpenStudy (anonymous):

\[\frac{x^2+8x+15}{x-4}\times \frac{x^2-16}{2x+6}\] \[\large=\frac{(x+5)(x+3)((x+4)(x-4)}{2(x-4)(x+3)}\] is a start

OpenStudy (anonymous):

it says "multiply" but it should actually say "factor and cancel"

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

get rid of any common factors you see top and bottom

OpenStudy (anonymous):

i did not explain how to factor you just kind of have to know that

OpenStudy (anonymous):

Thank you @satellite73 that helped alot

OpenStudy (anonymous):

i.e. you have to know that \(x^2-16=(x+4)(x-4)\) for example yw